# Bounded geometry for PCF-special subvarieties

### v1

@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
We now return to our proof. We know from Lemma 4.1 that zooming in at any exponential rate, the graphs of iterates of $c_{r+1}$ are converging to the graph of a constant function as we zoom in. In particular, recalling the definition of the bifurcation current associated to $c_{r+1}$, we can work in homogeneous coordinates the the parameter $s$ on a neighborhood of $s_{\ell}\in X_{\lambda_{1}}$, where $s_{\ell}$ is the last coordinate of $x_{1}=(\lambda_{1},s_{1},\ldots,s_{\ell})$ from Proposition 3.2. As $f_{s_{\ell}}^{n}(c_{r+1}(s_{\ell}))$ converges to a parabolic point for $f_{s_{\ell}}$, we define
$U_{n}(s):=\log\|(F_{s})^{n}(\tilde{c}_{r+1}(s))\|,$
with a choice of homogeneous lifts $F_{s}$ and $\tilde{c}_{r+1}$ for $f_{s}$ and $c_{r+1}$ normalized so that $U_{n}(s_{\ell})\to 0$ as $n\to\infty$. Note that the sequence of functions $d^{-n}U_{n}$ converges uniformly to the potential of the bifurcation current $T_{c_{r+1}}|_{X_{\lambda_{1}}}$ on a neighborhood of $s_{\ell}\in X_{\lambda_{1}}$. Let $r_{0}$ be any biholomorphic map from $\mathbb{D}^{r}$ to a neighborhood of $s_{\ell}\in X_{\lambda_{1}}$, with $r_{0}(0)=s_{\ell}$. The map $r_{0}$ can be viewed as a change of coordinates for a neighborhood of $s_{\ell}$. From Lemma 4.1, we deduce that, for any $\rho>1$, the functions $U_{n}\circ r_{0}(y/\rho^{n})$ converge uniformly to 00 on a small neighborhood of $y=(0,\cdots,0)\in\mathbb{D}^{r}$. As a consequence, for any $\rho_{i}\in\mathbb{C}$ with $|\rho_{i}|>1$ , we pick some $\rho>1$ with $\rho<\min_{1\leq i\leq{r}}(|\rho_{i}|)$, the sequence of functions
$U_{n}\circ r_{0}(y_{1}/\rho_{1}^{n},\ldots,y_{r}/\rho_{r}^{n})=U_{n}\circ r_{0}(A_{n}(y)/\rho^{n})\to 0$
conveges uniformly as $A_{n}(y):=(y_{1}\cdot\rho^{n}/\rho_{1}^{n},\ldots,y_{r}\cdot\rho^{n}/\rho_{r}^{n})$ converges to zero uniformly for $n\to\infty$. It follows that $r_{n}^{*}U_{n}$ also converges uniformly to 0, where $r_{n}$ is the notation for the zooming map of [Ga2, Proposition 3.2], which is a change of local coordinates at $s_{\ell}\in X_{\lambda_{1}}$ and then zoom in at some exponential rate in each coordinate; see [BE, Section 3] for more details.
Recall the continuity properties of currents, for example in [Dem]. In particular, if a sequence $u_{n}$ of plurisubharmonic functions converges uniformly to 0, then $dd^{c}u_{n}\to 0$ weakly. And if we have sequences $u_{n}^{i}$, $i=1,\ldots,{r}$, converging uniformly to $u^{i}$, then
$dd^{c}u_{n}^{1}\wedge\cdots\wedge dd^{c}u_{n}^{r}\to dd^{c}u^{1}\wedge\cdots\wedge dd^{c}u^{r}$
weakly. It follows that, by replacing just one of the prerepelling critical points (in our case that is $c_{1}$) with the critical point attracted to the parabolic cycle (i.e., $c_{r+1}$), the limit of the zooming process at the parameter $s_{\ell}\in X_{\lambda_{1}}$ for the new measure must be equal to 0. This completes the proof of Theorem 1.7.
@@
The proof would simplify if we knew that the $(r+1)$-th critical point $c_{r+1}$ were preperiodic to a parabolic cycle, instead of only in the basin of that cycle. We could employ ideas that appeared in [GNY, Proposition 5.2] (written for measures in the dynamical plane) or [FG3, Proposition 5.12] (for $1$-parameter families of polynomials and marked points). But, unfortunately, the critical point $c_{r+1}$ is sometimes the only “free” critical point and must therefore have an infinite orbit in the presence of a parabolic cycle.
@@
An elementary, but important, geometric observation is the following [GGK, Lemma 4.3]: If $\Sigma$ is a subset of a complex projective algebraic variety $X$, equipped with an ample line bundle $L$, and if the product $\Sigma^{M}$ lies in a proper subvariety $Y\subset X^{M}$ for some integer $M\geq 1$, then $\Sigma$ must be contained in a proper subvariety $Z$ of $X$, where the number of components of $Z$ and the degree of each are bounded by constants depending only on $M$, $\dim X$, the degree of $X$, and the sum of the degrees of the components of $Y$. Here, we measure degree in $X^{M}$ with respect to the ample line bundle $p_{1}^{*}L\otimes\cdots\otimes p_{M}^{*}L$ on $X^{M}$, where $p_{i}:X^{M}\to X$ is projection to the $i$-th factor.

### v2

@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
We now return to our proof. Abusing the notation slightly, we write $\{f_{s}\}_{s\in X_{\lambda_{1}}}$ to denote the restriction of $f_{\ell}$ on $X_{\lambda_{1}}$, $c_{r+1}$ for the restriction of $c_{r+1}$ to $X_{\lambda_{1}}$, and $T_{c_{r+1}}$ its associated bifurcation current on $X_{\lambda_{1}}$. We know from Lemma 4.1 that, zooming into $s_{\ell}$ in $X_{\lambda_{1}}$ at any exponential rate, the graphs of iterates of $c_{r+1}$ are converging to the graph of a constant function.
Let $p_{0}\in\mathbb{P}^{1}$ be the parabolic point of $f_{s_{\ell}}$; replacing $f_{s_{\ell}}$ with an iterate, we have $f_{s_{\ell}}^{n}(c_{r+1}(s_{\ell}))\to p_{0}$ as $n\to\infty$. After a change of coordinates if necessary, we may assume that $c_{r+1}(s_{\ell})$ and $p_{0}$ are not $\infty$. Recalling the definition of the bifurcation current associated to $c_{r+1}$, we can work in homogeneous coordinates on $\mathbb{P}^{1}$ over a neighborhood of $s_{\ell}$ in $X_{\lambda_{1}}$. We define
$G_{s}(X,Y):=\lim_{i\to\infty}\frac{1}{d^{i}}\log\|F_{s}^{i}(X,Y)\|,$
with a choice of homogeneous lift $F_{s}$ for $f_{s}$ normalized so that $G_{s_{\ell}}(p_{0},1)=0$. Note that $G_{s}(c_{r+1}(s),1)$ is a potential of the bifurcation current $T_{c_{r+1}}$ on a neighborhood of $s_{\ell}\in X_{\lambda_{1}}$. As long as $f^{n}_{s}(c_{r+1}(s))$ is not $\infty$, which is always true when $s$ is in $\mathbb{D}(s_{\ell},\delta/\rho^{n})$ for small $\delta>0$, any $\rho>1$ and big $n$ by Lemma 4.1, we define
$U_{n}(s):=G_{s}(f^{n}_{s}(c_{r+1}(s)),1)$
so that
$dd^{c}U_{n}(s)=d^{n}\cdot dd^{c}G_{s}(c_{r+1}(s),1)=d^{n}\,T_{c_{r+1}}$
where defined, as $G_{s}(F^{n}_{s}(c_{r+1}(s),1))=d^{n}\cdot G_{s}(c_{r+1}(s),1)$ differs from $U_{n}(s)$ by a harmonic function.
Let $r_{0}$ be any biholomorphic map from $\mathbb{D}^{r}$ to a neighborhood of $s_{\ell}\in X_{\lambda_{1}}$, with $r_{0}(0)=s_{\ell}$. From Lemma 4.1 and the choice of our lift, we deduce that, for any $\rho>1$, the functions $U_{n}\circ r_{0}(y/\rho^{n})$ converge uniformly to $G_{s_{\ell}}(p_{0},1)=0$ on a small neighborhood of $(0,\ldots,0)\in\mathbb{D}^{r}$. As a consequence, for any $\rho_{i}\in\mathbb{C}$ with $|\rho_{i}|>1$ , we can pick some $\rho>1$ with $\rho<\min_{1\leq i\leq{r}}(|\rho_{i}|)$ so that the sequence of functions
(4.10) $U_{n}\circ r_{0}(y_{1}/\rho_{1}^{n},\ldots,y_{r}/\rho_{r}^{n})=U_{n}\circ r_{0}(A_{n}(y)/\rho^{n})\to 0$
uniformly, as $A_{n}(y):=(y_{1}\cdot\rho^{n}/\rho_{1}^{n},\ldots,y_{r}\cdot\rho^{n}/\rho_{r}^{n})$ converges to zero uniformly for $n\to\infty$.
Following the notation of [Ga2, Proposition 3.2], we now let $r_{n}:\mathbb{D}^{r}\to X_{\lambda_{1}}$ be the sequence of maps that define the zooms near the parameter $s_{\ell}$ for the measure $\mu^{r+1}_{\lambda_{1}}$, as described before Lemma 4.1. It follows from (4.10) that the functions $U_{n}\circ r_{n}$ also converge uniformly to 0. We conclude that $d^{n}\,(r_{n})^{*}T_{c_{r+1}}\to 0$ weakly, and therefore
$d^{nr}\,(r_{n})^{*}\mu^{1}_{\lambda_{1}}\to 0.$
We have thus reached our desired contradiction, and this completes the proof of Theorem 1.8.
@@
The proof would simplify if we knew that the $(r+1)$-th critical point $c_{r+1}$ were preperiodic to a parabolic cycle, instead of only in the basin of that cycle. We could employ ideas that appeared in [GNY, Proposition 5.2] (written for measures in the dynamical plane) or [FG3, Proposition 5.12] (for $1$-parameter families of polynomials and marked points). But the critical point $c_{r+1}$ is sometimes the only “free” critical point and must therefore have an infinite orbit in the presence of a parabolic cycle.
@@
An elementary, but important, geometric observation is the following [GGK, Lemma 4.3]: If $\Sigma$ is a subset of a complex projective algebraic variety $X$, equipped with an ample line bundle $L$, and if the product $\Sigma^{M}$ lies in a proper subvariety $Y\subset X^{M}$ for some integer $M\geq 1$, then $\Sigma$ must be contained in a proper subvariety $Z$ of $X$, where the number of components of $Z$ and the degree of each are bounded by constants depending only on $M$, $\dim X$, the degree of $X$, and the sum of the degrees of the components of $Y$. Here, we measure degree in $X^{M}$ with respect to the ample line bundle $p_{1}^{*}L\otimes\cdots\otimes p_{M}^{*}L$ on $X^{M}$, where $p_{i}:X^{M}\to X$ is the projection to the $i$-th factor.